The sum of two consecutive integers is 527. What are the two numbers?
Answer by checkley77(12844)

x+(x+1)=527
2x=527-1
2x=526
x=526/2
x=263 ans. for the smaller integer.
263+1=264 ans. for the larger integer.
Proof:
263+264=527
527=527